OpenStudy (gavin39):

An airplane takes 4 hours to travel a distance of 2400 miles with the wind. The return trip takes 5 hours against the wind. Find the speed of the plane in still air and the speed of the wind.

OpenStudy (gavin39):

Last question on my quiz if anyone can help real fast

OpenStudy (gavin39):

homework***

OpenStudy (ben00):

try dividing, divide 2400 by 400.

OpenStudy (danjs):

Here are vectors for the velocity of the plane and wind 

OpenStudy (gavin39):

So it would be 600 mph for the plane? I thought it would be 600, but kept confusing myself. Especially with the wind speed

OpenStudy (gavin39):

So for the wind it would be 600/5

OpenStudy (gavin39):

Im assuming

OpenStudy (danjs):

overall velocity on the way there V = P + W Overall on the way back v = P - W

OpenStudy (danjs):

(2400/4 ) mi/h = P + W (2400/5) mi/h = P - W solve that

OpenStudy (ben00):

how fast is the wind?

OpenStudy (gavin39):

120 I believe?

OpenStudy (ben00):

alright so plane speed is 600 and wind speed is 120.

OpenStudy (gavin39):

Yeah I'm fairly sure 120

OpenStudy (gavin39):

Awesome. Thanks guys!

OpenStudy (ben00):

give medal to us then.

OpenStudy (danjs):

Wind W and PLane P 600 = P + W 480 = P - W Solving that gives, P=540, W=60